Companies are interested in the demographic of those who listen to the radio programs they sponsor. A radio station has determined that only 20% of listeners phoning into a morning talkback program are male. During a particular week, 200 calls are received by this program.
a) What is the probability that at least 50 of these 200 callers are male?
b) What is the probability that more than half of these 200 callers are female?
c) There is a 30% chance that the number of male callers among the 200 total callers does
not exceed what?

Answer by ewatrrr(24785)


Hi
Binomial Distribution:  p = .20 male  q = .80 female   n = 200
For comparison: .20*200 = 40
 P(x ≥ 50m) = 1 - P(x ≤ 49m) = 1 - .9506 = .0474

 P(x > 100f) = 1 - P(x ≤ 100f)  = basically 1 or 100%

200*.30 = 60 
 p(x = 60) = .0002  0r .02%

Wish You the Best in your Studies.